### v2

# ARFlow: Human Action-Reaction Flow Matching with Physical Guidance

Abstract
Human action-reaction synthesis, a fundamental challenge in modeling causal human interactions, plays a critical role in applications ranging from virtual reality to social robotics. While diffusion-based models have demonstrated promising performance, they exhibit two key limitations for interaction synthesis: reliance on complex noise-to-reaction generators with intricate conditional mechanisms, and frequent physical violations in generated motions. To address these issues, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that establishes direct action-to-reaction mappings, eliminating the need for complex conditional mechanisms. Our approach introduces two key innovations: an $\mathbf{x}_{1}$-prediction method that directly outputs human motions instead of velocity fields, enabling explicit constraint enforcement; and a training-free, gradient-based physical guidance mechanism that effectively prevents body penetration artifacts during sampling. Extensive experiments on NTU120 and Chi3D datasets demonstrate that ARFlow not only outperforms existing methods in terms of Fréchet Inception Distance and motion diversity but also significantly reduces body collisions, as measured by our new Intersection Volume and Intersection Frequency metrics.
Figure 1: Our proposed Human Action-Reaction Flow (ARFlow). We directly establish a mapping between the action and reaction distribution and our sampling process is further guided by our physical constraint guidance. The change of colors represents the variation of the $h$-frame reaction with sampling timestep $t_{n}$.
$\dagger$\$$\dagger$\$footnotetext: Corresponding authors.
1 Introduction
Human action-reaction synthesis [56, 69, 10] has emerged as a pivotal research direction in computer vision [67, 37, 55, 26, 57, 64]. This task aims to generate physically plausible human reactions responding to observed actions, with critical applications in virtual reality, human-robot interaction, and character animation. Unlike single-human motion generation [18, 59, 8], reactors must infer responses without observing future actor motions, creating unique modeling challenges.
While recent diffusion methods [59, 69] show promise in motion generation, they face two key limitations in the modeling of action-reaction interactions. First, existing approaches  [59, 69] indirectly model responses using noise-to-reaction generators with intricate conditional mechanisms like treating action information as a condition to guide the generation process. This not only complicates the training process but also fails to directly capture the inherent causal relationship between actions and reactions. Second, frequent physical violations like body penetration between characters occur due to neglected physical constraints. While such issues are absent in single-human scenarios, they become critical in human interaction applications [24]. This poses a significant barrier to real-world applications such as virtual reality and human-robot interaction, where even minor physical inaccuracies are intolerable [50, 24].
To address these challenges, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that fundamentally resolves these limitations. Unlike diffusion models constrained by noise-data mappings, flow matching [38, 42, 1, 45] naturally models paired distributions through linear interpolation between endpoints (See Fig. 1), making it inherently suitable for action-reaction synthesis. ARFlow establishes direct pathways between action and reaction distributions, enabling simpler training and faster inference compared to diffusion-based methods. Instead of traditional vector field prediction ($\mathbf{v}$-prediction) in standard flow matching [38, 25], ARFlow directly outputs clean body poses ($\mathbf{x}_{1}$-prediction). This innovation allows direct application of geometric constraints during training and better alignment with motion generation objectives. We further develop a specialized sampling algorithm optimized for $\mathbf{x}_{1}$-prediction in human motion generation. To eliminate unrealistic body collisions between characters, we introduce a penetration-aware sampling process that maintains physical plausibility through gradient guidance without compromising motion quality. Additionally, we introduce two new physics-aware evaluation metrics: Intersection Volume (IV) and Intersection Frequency (IF). Our main contributions are threefold:
  • We propose ARFlow, the first flow matching architecture that creates direct pathways between human action and reaction distributions, eliminating the need for complex conditional mechanisms prevalent in diffusion approaches.
  • We propose a new $\mathbf{x}_{1}$-prediction method under flow matching specifically designed to integrate physical constraints during both training and sampling. This combined approach successfully prevents body collisions between characters while preserving natural movement.
  • We introduce two new evaluation metrics (IV, IF) for interaction quality assessment. Experiments demonstrate state-of-the-art performance in Fréchet Inception Distance (FID), MultiModality (Multimod.), IV and IF on NTU120 and Chi3D datasets.
Figure 2: Pipeline of ARFlow. (a) For a sampled timestep $t$, we linearly interpolate a coupled action-reaction pair as Eq. 4 to produce the intermediate state $x^{1:H}_{t}$, which is then turns into a d-dimensional latent feature through a linear layer. We use Transformer Decoder Units to directly predict clean reaction motions. (b) After training the networks in (a), our ARFlow uses them for $\mathbf{x}_{1}$-$prediction$ based sampling. The sampling process is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
2 Related Work
2.1 Human Motion Generation
Human motion synthesis aims to generate diverse and realistic human-like motion conditioned on different guidances, including text [76, 19, 31, 14], music [80, 35], speech [20, 4], sparse signals [15, 58], action labels [18, 47, 59, 6, 8], or unconditioned [48, 59, 71]. Recently, many diffusion-based motion generation models have been proposed [75, 59, 14, 8, 66, 60] and demonstrate better quality compared to alternative models such as VAE [32, 18, 47, 6], flow-based models [51, 3] or GANs [71, 67]. Alternatively, motion can be regarded as a new form of language and embedded into the language model framework [74, 27].
Meanwhile, the exploration of guiding the sampling process of diffusion models [12, 72] has been a key area in motion diffusion models, PhysDiff [29] proposes a physics-guided motion diffusion model, which incorporates physical constraints in a physics simulator into the diffusion process. GMD [29] presents methods to enable spatial guidance without retraining the model for a new task. DNO [29] proposes a motion editing and control approach by optimizing the diffusion latent noise of an existing pre-trained model.
2.2 Human Action-Reaction Synthesis
Different from human-human interaction [67, 37, 55, 26, 57, 64], human action-reaction synthesis is causal and asymmetric [69, 40, 68]. To address this task, researchers have leveraged large language models [54, 56, 28], VAE-based methods [10, 43, 44]. However, these methods cannot capture fine-grained representations and ensure diversity, and diffusion-based methods [36, 57] are limited to the “offline” and “constrained” setting of human reaction generation, failing to generate instant and intention agnostic reactions. More recently, ReGenNet [69] introduce a diffusion-based transformer decoder framework and treat action sequence as conditional signal for online reaction generation. However, it often produce physically-implausible inter-penetrations between the actor and reactor since they disregard physical constraints in the generative process. Our method addresses this problem by guiding the sampling process with physical constraint guidance.
2.3 Flow Matching
Flow Matching [38, 42, 1, 45] has emerged as an efficient alternative to diffusion models, offering linear generation trajectories through ODE solvers. This paradigm enables simplified training and accelerated inference [16, 39], with successful applications spanning images [38, 16], audio [34], video [5], and point clouds [65]. In motion generation, MotionFlow [25] demonstrates comparable performance to diffusion models with faster sampling. Notably, Flow Matching inherently models transitions between arbitrary distributions through transport maps, making it particularly suitable for paired data modeling. Despite these advantages, its potential for action-reaction synthesis remains unexplored. Our work bridges this gap by establishing direct action-to-reaction mappings without complex conditional mechanisms.
3 Method
In the setting of human action-reaction synthesis, our primary goal is to generate the reaction $\mathbf{x}_{1}=\{x_{1}^{i}\}_{i=1}^{H}$ conditioned on an arbitrary action $\mathbf{x}_{0}=\{x_{0}^{i}\}_{i=1}^{H}$ of length $H$. The condition $\mathbf{c}$ can be action $\mathbf{x}_{0}$, or it can be a signal such as an action label, text, audio to instruct the interaction, which is optional for intention-agnostic scenarios. We utilize SMPL-X [46] human model to represent the human motion sequence as [69] to improve the modeling of human-human interactions. Thus, the reaction can be represented as $x^{i}_{1}=\left[\bm{\theta}^{x_{1}}_{i},\bm{q}^{x_{1}}_{i},\bm{\gamma}^{x_{1}}_{i}\right]$ where $\bm{\theta}^{x_{1}}_{i}\in\mathbb{R}^{3K}$, $\bm{q}^{x_{1}}_{i}\in\mathbb{R}^{3}$, $\bm{\gamma}^{x_{1}}_{i}\in\mathbb{R}^{3}$ are the pose parameters, the global orientation, and the root translation of the person, respectively. Total number $K$ of body joints, including the jaw, eyeballs, and fingers, is 54. The main pipeline of our ARFlow model is provided in Fig. 2. In this section, we first introduce the Human Action-Reaction Flow Matching in Sec. 3.1. Then, we present our training-free physical constraint guidance method to address the issue of physically implausible human-human inter-penetrations in Sec. 3.2.
3.1 Human Action-Reaction Flow Matching
Flow Matching Overview.
Given a set of samples from an unknown data distribution $q(\mathbf{x})$, the goal of flow maching is to learn a flow that transforms a prior distribution $p_{0}(\mathbf{x})$ towards a target data distribution $p_{1}(\mathbf{x}){\approx}q(\mathbf{x})$ along the probability path $p_{t}(\mathbf{x})$. The time-dependent flow $\phi_{t}(\mathbf{x})$ is defined by a vector field $\mathbf{v}(\mathbf{x},t):\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ which establishes the flow through a neural ODE:
$\frac{d}{dt}\phi_{t}(\mathbf{x})=\mathbf{v}(\phi_{t}(\mathbf{x}),t),\quad\quad\phi_{0}(\mathbf{x})$ $=\mathbf{x}~{}.$ (1)
Given a predefined probability path $p_{t}(\mathbf{x})$ and a corresponding vector field $\mathbf{u}_{t}(\mathbf{x})$, one can regress the vector field $\mathbf{u}_{t}(\mathbf{x})$ with a neural network $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ parameterized by $\theta$, and the Flow Maching (FM) objective is as follows:
$\min_{\theta}\mathbb{E}_{t,p_{t}(\mathbf{x})}\|\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-\mathbf{u}_{t}(\mathbf{x})\|^{2}.$ (2)
By defining the conditional probability path as a linear interpolation between $p_{0}$ and $p_{1}$, the intermediate process becomes: $\mathbf{x}_{t}=t\mathbf{x}_{1}+[1-(1-\sigma_{\text{min}})t]\mathbf{x}_{0}$, where $\sigma_{\text{min}}>0$ is a small amount of noise. Both training and sampling are simplified by fitting a linear trajectory in contrast to diffusion paths. When extra condition signals $\mathbf{c}$ are required, they can be directly incorporated into the vector field estimator $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ as $\mathbf{v}(\mathbf{x}_{t},t,\mathbf{c})$. Therefore, the training objective is as follows:
$\min_{\theta}\mathbb{E}_{t,p(\mathbf{x}_{0}),q(\mathbf{x}_{1})}\Big{\|}\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)-\Bigl{(}\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0}\Bigr{)}\Big{\|}^{2}.$ (3)
Action-Reaction Flow Matching.
Different from previous diffusion-based methods[69, 59, 36, 15] that rely on cumbersome conditional mechanisms, we adopt Flow Matching to directly construct a mapping from action distribution to reaction distribution (See Fig. 1). Specially, we build a generative model $f$ parametrized by $\theta$ to synthesize the reaction $\mathbf{x}_{1}=f_{\theta}(\mathbf{x}_{0},\mathbf{c})$, given action $\mathbf{x}_{0}$, instead of $\mathbf{x}_{1}=f_{\theta}(\mathbf{z},\mathbf{c})$ in diffusion, given a sampled Gaussian noise vector $\mathbf{z}$. Given the reaction $\mathbf{x}_{1}$ sampled from the reaction distribution and the coupled action $\mathbf{x}_{0}$ from the action distribution, the intermediate process can be written as
${\mathbf{x}_{t}=t\mathbf{x}_{1}+[1-(1-\sigma_{\text{min}})t]\mathbf{x}_{0},}$ (4)
where $t$ is the timestep, $\sigma_{\text{min}}>0$ is a small amount of noise. In our setting, the generative process is conditionally formulated as $p(\mathbf{x}_{t_{n+1}}|\mathbf{x}_{t_{n}},c)$. We use a neural network $G$ to directly predict the clean body poses, i.e., $\hat{\mathbf{x}}_{1}=G_{\theta}(\mathbf{x}_{t},t,c)$, instead of predicting vector fields in previous works [25, 38]. This strategy is both straightforward and effective, since many geometric losses directly act on the predicted $\hat{\mathbf{x}}_{1}$. We compared and analyzed the results of predicting vector fields ($\mathbf{v}$-prediction) and clean body poses ($\mathbf{x}_{1}$-prediction) in Sec. 4.4. Note that $\mathbf{x}_{1}$ in flow maching usually corresponds to $\mathbf{x}_{0}$ in previous literature on diffusion models. Depending on the specific application, $G$ can be implemented by Transformers [62] or MLP networks. The training objective of our flow model is as follows:
${\mathcal{L}_{\text{fm}}=\mathbb{E}_{\mathbf{x}_{1}\sim q(\mathbf{x}_{1}),\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),t\sim[0,1]}[\|\mathbf{x}_{1}-G_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}].}$ (5)
Following [69], we employ explicit interaction losses to evaluate the relative distances of body pose $\bm{\theta}(\mathbf{x}_{1},\mathbf{x}_{0})$, orientation $\bm{q}(\mathbf{x}_{1},\mathbf{x}_{0})$ and translation $\bm{\gamma}(\mathbf{x}_{1},\mathbf{x}_{0})$ between the actor and reactor. We use a forward kinematic function to transforms the rotation pose into joint positions for calculating $\bm{\theta}(\mathbf{x}_{1},\mathbf{x}_{0})$ , and converts the rotation poses to rotation matrices for calculating $\bm{q}(\mathbf{x}_{1},\mathbf{x}_{0})$. The interaction loss is defined as
$\mathcal{L}_{\text{inter}}=\frac{1}{H}\biggl{(}\|\bm{\theta}(\mathbf{x}_{1},\mathbf{x}_{0})-\bm{\theta}(\hat{\mathbf{x}}_{1},\mathbf{x}_{0})\|_{2}^{2}$ (6)
$+\|\bm{q}(\mathbf{x}_{1},\mathbf{x}_{0})-\bm{q}(\hat{\mathbf{x}}_{1},\mathbf{x}_{0})\|_{2}^{2}+\|\bm{\gamma}(\mathbf{x}_{1},\mathbf{x}_{0})-\bm{\gamma}(\hat{\mathbf{x}}_{1},\mathbf{x}_{0})\|_{2}^{2}\biggr{)}.$
Our overall training loss is $\mathcal{L}_{\text{all}}=\mathcal{L}_{\text{fm}}+\lambda_{\text{inter}}\cdot\mathcal{L}_{\text{inter}}$, and $\lambda_{\text{inter}}$ is the loss weight.
Sampling based on $\mathbf{x}_{1}$-prediction.
Since our neural network outputs $\hat{\mathbf{x}}_{1}$, we require to construct an equivalent relationship between the neural network’s predictions of $\mathbf{v}$ and $\mathbf{x}_{1}$. The equivalent form of parameterization Equation 21 derived from our appendix is as follows:
${\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t},}$ (7)
Then, our sampling based on $\mathbf{x}_{1}$-prediction can be achieved by first sampling $\mathbf{x}_{0}$ and then solving Equation 1 employing an ODE solver [52, 33, 2] through our trained neural network $\hat{\mathbf{x}}_{1}=G_{\theta}(\mathbf{x}_{t},t,c)$. We use the Euler ODE solver and discretization process involves dividing the procedure into $N$ steps, leading to the following formulation:
$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}_{t_{n}}+(t_{n+1}-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (8)
where the integer time step $t_{1}=0<t_{2}<\cdots<t_{N}=1$. By using equivalent form of parameterization Equation 7, we finally obtain our flow maching sampling formulation based on $\mathbf{x}_{1}$-prediction:
$\mathbf{x}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1},$ (9)
which is more suitable for human motion generation. Detailed derivation is provided in the supplementary.
3.2 Physical Constraint Guidance
There exist many physically implausible inter-penetrations between the actor and reactor in the generated results of current diffusion-based methods [69, 59, 15], owing to the stochastic sampling process that disregards physical constraints.
To address the issue, we propose a training-free physical constraint guidance method that calculates the penetration gradients $\nabla\mathcal{L}_{\text{pene}}$ that pulls the reactor in the opposite direction to the actor. Then these gradients are employed to guide the sampling process at each timestep to ensure the suppression of penetration. Specifically, We use $\psi^{h}_{i}(\mathbf{x})$ to represent the position of the $i$-th joint in the $h$-th frame of the generated reaction motion $\mathbf{x}$. Inspired by [30, 36], we calculate the signed distance function (SDF) between the actor and the reactor. Thus, our penetration loss functon is defined as:
$\mathcal{L}_{\text{pene}}(\mathbf{x})\vcentcolon=\sum_{i,h}-\mathrm{min}\Bigl{(}\mathrm{SDF}(\psi^{h}_{i}(\mathbf{x})),\zeta\Bigr{)}$ (10)
where $\zeta$ defines the safe distance between the actor and the reactor, beyond which the gradient becomes zero, and $\mathrm{SDF}$ is the signed distance function for an actor in the $h$-th frame, which dynamically changes across frames.
Flow maching sampling with physical constraint guidance. At each sampling step, the gradient of $\nabla\mathcal{L}_{\text{pene}}$ is utilized to guide the generative process. In contrast to the standard process for motion synthesis in Equation 8, the sampling process with physical constraint guidance is formalized as:
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\mathbf{x}_{t_{n}}+(t_{n+1}-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (11)
$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (12)
where $\mathbf{x}^{\prime}_{t_{n+1}}$ denoted the intermediate results of standard sampling in Equation 8. However, standard sampling based on $\mathbf{v}$-prediction often requires to estimate $\hat{\mathbf{x}}_{1}$:
$\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (13)
from current state $\mathbf{x}_{t_{n}}$ using an equivalent form of parameterization in Equation 7 to input into the signed distance function $\mathcal{L}_{\text{pene}}(\mathbf{x})$. In comparison, the sampling process based on $\mathbf{x}_{1}$-prediction with physical constraint guidance is as follows:
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1},$ (14)
$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (15)
where $\hat{x}_{1}$ is the clean body poses predicted by a neural network $G_{\theta}$, which is more compatible with our physical constraint guidance method. The complete Euler sampling with physical constraint guidance for $\mathbf{x}_{1}$-prediction or $\mathbf{v}$-prediction is shown in Algorithm 1 and Algorithm 2 respectively. In practice, we use $\mathbf{x}_{1}$-prediction for its better performance. Under iterative sampling and physical constraint guidance, our method can generate more realistic and physically-plausible reaction motions.
Our guidance method is actually a refined fine-tuning, which may be not suitable for training phases. In addition, the loss during the training phase mainly measures the difference between generated results and ground truth, while our guidance during the inference phase can provide more flexible guidance based on the quality of the generated results.
Metrics. To qualitatively measure the degree of penetration, we introduced two metrics: 1) Intersection Volume (IV). Penetrate in [73, 21] just measures ground penetration which is not suitable for measuring the degree of penetration between humans. Interpenetration in [44] can only be computed as rigid bodies in the physics simulation. Inspired by Solid Intersection Volume (IV) [78, 41] in hand-object interaction, we measure human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both. Intersection Volume (IV) is defined as
$IV=\frac{1}{H\cdot N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}\sum_{h=1}^{H}V^{h}_{\text{pene}},$ (16)
where $V^{h}_{\text{pene}}$ represents the intersection volume of frame $h$ and $N_{\text{total}}$ denotes the total number of samples.
2) Intersection Frequency (IF). However, solely considering this metric can be misleading since it does not account for the case where there is no contact between the reactor and the reactor at all. Inspired by Contact Frequency in [36, 54], we introduce IF to measure the frequency of inter-penetration, i.e.
$IF=\frac{f_{\text{pene}}}{F_{\text{total}}}\cdot 100\%,$ (17)
where $f_{\text{pene}}$ represents the number of frames with inter-penetrations and $F_{\text{total}}$ denotes the total number of frames.
Algorithm 1 Euler Sampling algorithm with physical constraint guidance. ($\mathbf{x}_{1}$-prediction)
1:  Input: $\mathcal{L}_{\text{pene}}$ the loss function ; $G$ and $\theta$ the clean body poses predictor with pretrained parameters
2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the scale factor to control the strength of guidance
3:  Sample $\mathbf{x}_{0}$ from the action distribution
4:  for $n=1,2,...,N-1$ do
5:     $\hat{\mathbf{x}}_{1}\leftarrow G_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
6:     # Flow maching $\mathbf{x}_{1}$-prediction sampling (Eq. 14)
7:     $\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1}$
8:     # Physical constraint guidance (Eq. 15)
9:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
10:  end for
11:  Return: The reaction motion after guidance $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
4 Experiments
Figure 3: Qualitative comparisons of human action-reaction synthesis results. Blue for actors and Green for reactors.
Our experiment setting of human action-reaction synthesis is online, unconstrained as in  [69] for its significant potential for practical applications. Online represents real-time reaction generation where future motions of the actor are not visible to the reactor, and the opposite is offline to relax the synchronicity. Unconstrained means that the intention of the actor is invisible to the reactor. To demonstrate the universality of our flow matching and physical guidance methods, we also conducted offline setting experiments.
[…]
Real $0.09^{\pm 0.00}$ $0.867^{\pm 0.0002}$ $13.06^{\pm 0.09}$ $25.03^{\pm 0.23}$
cVAE [32] $70.10^{\pm 3.42}$ $0.724^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.40^{\pm 0.26}$
AGRoL [15] $44.94^{\pm 2.46}$ $0.680^{\pm 0.0001}$ $12.51^{\pm 0.09}$ $19.73^{\pm 0.17}$
MDM [59] $54.54^{\pm 3.94}$ $0.704^{\pm 0.0003}$ $11.98^{\pm 0.07}$ $19.45^{\pm 0.20}$
MDM-GRU [59] $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
ReGenNet $\underline{11.00^{\pm 0.74}}$ $\mathbf{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
ARFlow $\mathbf{7.89^{\pm 0.18}}$ $\underline{0.743^{\pm 0.0002}}$ $\underline{13.60^{\pm 0.10}}$ $\mathbf{24.11^{\pm 0.13}}$
Table 1: Comparison to state-of-the-art on the online, unconstrained setting on NTU120-AS. $\rightarrow$ denotes that the result closer to Real is better, and $\pm$ represents 95% confidence interval. We highlight the best result in Bold and the second best in underline.
[…]
Real $0.75^{\pm 0.18}$ $0.691^{\pm 0.0093}$ $7.15^{\pm 1.27}$ $12.94^{\pm 0.96}$
cVAE [32] $17.33^{\pm 17.14}$ $0.552^{\pm 0.0024}$ $8.20^{\pm 0.57}$ $11.44^{\pm 0.35}$
AGRoL [15] $64.83^{\pm 277.8}$ $\underline{0.644^{\pm 0.0039}}$ $\mathbf{7.00^{\pm 0.95}}$ $11.33^{\pm 0.65}$
MDM [59] $18.40^{\pm 7.95}$ $\mathbf{0.647^{\pm 0.0035}}$ $5.89^{\pm 0.33}$ $10.96^{\pm 0.27}$
MDM-GRU [59] $18.63^{\pm 25.87}$ $0.574^{\pm 0.0046}$ $6.20^{\pm 0.24}$ $10.49^{\pm 0.32}$
ReGenNet $\underline{13.76^{\pm 4.78}}$ $0.601^{\pm 0.0040}$ $6.35^{\pm 0.24}$ $\underline{12.02^{\pm 0.33}}$
ARFlow $\mathbf{10.92^{\pm 3.70}}$ $0.600^{\pm 0.0040}$ $\underline{6.68^{\pm 0.25}}$ $\mathbf{12.74^{\pm 0.17}}$
Table 2: Comparison to state-of-the-art on the online, unconstrained setting on Chi3D-AS. $\rightarrow$ denotes that the result closer to Real is better, and $\pm$ represents 95% confidence interval. We highlight the best result in Bold and the second best in underline.
[…]
Real $0.09^{\pm 0.00}$ $0.867^{\pm 0.0002}$ $13.06^{\pm 0.09}$ $25.03^{\pm 0.23}$
cVAE [32] $74.73^{\pm 4.86}$ $0.760^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.40^{\pm 0.26}$
AGRoL [15] $16.55^{\pm 1.41}$ $0.716^{\pm 0.0002}$ $13.84^{\pm 0.10}$ $21.73^{\pm 0.20}$
MDM [59] $7.49^{\pm 0.62}$ $\mathbf{0.775^{\pm 0.0003}}$ $\underline{13.67^{\pm 0.18}}$ $24.14^{\pm 0.29}$
MDM-GRU [59] $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
ReGenNet $\underline{6.19^{\pm 0.33}}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\underline{25.21^{\pm 0.34}}$
ARFlow $\mathbf{5.00^{\pm 0.17}}$ $\underline{0.772^{\pm 0.0002}}$ $13.84^{\pm 0.09}$ $\mathbf{25.10^{\pm 0.17}}$
Table 3: Results on the offline, unconstrained setting on NTU120-AS. We highlight the best result in Bold and the second best in underline.
Table 4: Human action-reaction synthesis with physical constraint guidance on NTU120-AS. Bold indicates the best result.
FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
Real 0.09 38.38% 3.33
       Offline
MDM 22.04 29.77% 2.50
ReGenNet 6.19 35.77% 2.85
ARFlow 5.00 23.91% 2.01
MDM ${}^{\text{guidance}}$ 20.42 25.58% 2.10
ReGenNet ${}^{\text{guidance}}$ 7.69 25.10% 1.92
ARFlow ${}^{\text{guidance}}$ 5.19 15.82% 1.31
       Online
MDM 51.11 51.87% 4.22
ReGenNet 11.00 31.92% 2.63
ARFlow 7.89 17.44% 1.55
MDM ${}^{\text{guidance}}$ 40.03 44.84% 3.67
ReGenNet ${}^{\text{guidance}}$ 11.63 17.11% 1.37
ARFlow ${}^{\text{guidance}}$ 8.75 8.56% 0.76
[…]
Real $0.085^{\pm 0.0003}$ $0.867^{\pm 0.0002}$ $13.063^{\pm 0.0908}$ $25.032^{\pm 0.2332}$
Prediction 1) $x_{1}$ $\mathbf{7.894^{\pm 0.1814}}$ $\mathbf{0.743^{\pm 0.0002}}$ $\mathbf{13.599^{\pm 0.1005}}$ $\mathbf{24.105^{\pm 0.1310}}$
2) $v$ $14.726^{\pm 0.2143}$ $0.743^{\pm 0.0002}$ $14.154^{\pm 0.0923}$ $23.329^{\pm 0.1125}$
Guidance w. $\mathcal{L}_{\text{pene}}$ $8.751^{\pm 0.2047}$ $0.740^{\pm 0.0002}$ $13.613^{\pm 0.1004}$ $24.096^{\pm 0.1433}$
Timesteps 2 $15.965^{\pm 0.2728}$ $0.733^{\pm 0.0002}$ $13.740^{\pm 0.0896}$ $26.767^{\pm 0.1440}$ $0.055$
5 $\mathbf{7.894^{\pm 0.1814}}$ $0.743^{\pm 0.0002}$ $\mathbf{13.599^{\pm 0.1005}}$ $\mathbf{24.105^{\pm 0.1310}}$ $0.111$
10 $8.273^{\pm 0.3862}$ $0.721^{\pm 0.0002}$ $14.108^{\pm 0.0779}$ $22.995^{\pm 0.1274}$ $0.232$
100 $8.259^{\pm 0.3902}$ $0.747^{\pm 0.0002}$ $14.173^{\pm 0.1024}$ $23.619^{\pm 0.1214}$ $2.273$
Best ARFlow $\mathbf{7.894^{\pm 0.1814}}$ $0.743^{\pm 0.0002}$ $\mathbf{13.599^{\pm 0.1005}}$ $\mathbf{24.105^{\pm 0.1310}}$ $0.111$
Table 5: Ablation studies on the online, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
[…]
Real $0.085^{\pm 0.0003}$ $0.867^{\pm 0.0002}$ $13.063^{\pm 0.0908}$ $25.032^{\pm 0.2332}$
Prediction 1) $x_{1}$ $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$
2) $v$ $7.585^{\pm 0.1562}$ $0.757^{\pm 0.0002}$ $13.775^{\pm 0.0982}$ $24.200^{\pm 0.1355}$
Guidance w. $\mathcal{L}_{\text{pene}}$ $5.196^{\pm 0.1144}$ $0.753^{\pm 0.0002}$ $13.824^{\pm 0.0839}$ $25.164^{\pm 0.1630}$
Timesteps 2 $7.936^{\pm 0.1581}$ $0.759^{\pm 0.0002}$ $14.538^{\pm 0.1016}$ $25.904^{\pm 0.1754}$ $0.023$
5 $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$ $0.053$
10 $5.506^{\pm 0.1657}$ $0.744^{\pm 0.0002}$ $13.870^{\pm 0.0942}$ $24.732^{\pm 0.1533}$ $0.110$
100 $5.836^{\pm 0.3763}$ $0.748^{\pm 0.0002}$ $13.635^{\pm 0.0948}$ $24.058^{\pm 0.1371}$ $1.132$
Best ARFlow $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$ $0.053$
Table 6: Ablation studies on the offline, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
4.1 Experiment setup
Evaluation Metrics. 1) We adopt the following metrics to quantitatively evaluate results: Frechet Inception Distance (FID), Action Recognition Accuracy (Acc.), Diversity (Div.) and Multi-modality (Multimod.). For all these metrics widely used in previous human motion generation [18, 47, 59, 69], we use the action recognition model [70] to extract motion features for calculating these metrics as in [69]. We generate 1,000 reaction samples by sampling actor motions from test sets and evaluate each method 20 times using different random seeds to calculate the average with the 95% confidence interval as prior works [18, 47, 59, 69]. 2) Intersection Volume (IV) measures human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both. 3) Intersection Frequency (IF) measures the frequency of inter-penetration. More details about these metrics are provided in the supplementary.
Datasets. We evaluate our model on NTU120-AS and Chi3D-AS datasets with SMPL-X [46] body models and actor-reactor annotations as in [47].They contains 8118 and 373 human interaction sequences, respectively. “AS”[69] represents that they are an extended version of the original dataset [17, 40, 61], which adds annotations to distinguish actor-reactor order of each interaction sequence and SMPL-X body models for more detailed representations. We adopt the 6D rotation representation [79] in all our experiments.
4.2 Comparison to baselines
To evaluate the performance of our method, we adopt following baselines: 1) cVAE [32], commonly utilized in earlier generative models for human interactions; 2) MDM [59], the state-of-the-art diffusion-based method for human motion generation, and its variant MDM-GRU [59], which incorporates a GRU [9] backbone; 3) AGRoL [15], the current state-of-the-art method to generate full-body motions from sparse tracking signals, which adopts diffusion models with MLPs architectures; 4) ReGenNet [69], the state-of-the-art diffusion-based method for human action-reaction synthesis on online, unconstraint setting as ours. Results are taken from tables of ReGenNet [69] where all methods use 5-timestep sampling.
For the NTU120-AS dataset in Tab. 1 and Chi3D-AS dataset in Tab. 2, our proposed ARFlow notably outperforms baselines in terms of the FID metric, demonstrating that our method better models the mapping between the action and the reaction distribution. Our method achieves the best FID and multi-modality, second best for the action recognition accuracy and diversity on the NTU120-AS dataset and the best FID and multi-modality, second best for the diversity on the Chi3D-AS dataset. For a fair comparison, we use the pre-trained action recognition model in [69], so our action recognition accuracy is very close to the results of [69]. Given the restricted size of the Chi3D-AS test set, some fluctuations in the experimental results are to be expected.
As for the offline setting, we replace the Transformer decoder units equipped with attention masks with an 8-layer Transformer encoder architecture just like ReGenNet [69]. As demonstrated in Tab. 3, our model achieves superior performance on most of the metrics, highlighting the effectiveness and versatility of our method across various scenarios.
4.3 Physical Constraint Guidance
Qualitative results. In Fig. 3, we provide a qualitative comparison of reaction sequences generated by MDM, ReGenNet and our ARFlow with physical constraint guidance, for the same action sequences sampled from the test sets of NTU120-AS datasets. The visualization results demonstrate that both MDM and ReGenNet produce varying degrees of penetration between the actor and the reactor. In contrast, our method not only produces more physically plausible reactions, i.e., mitigate penetration problems, but also more responsive reactions, i.e., move backward when kicked, due to stronger modeling ability for causal relationship between actions and reactions. For more visualizations and videos of the generated human reactions, please refer to the supplementary materials.
Quantitative results. In Tab. 4, our ARFlow with physical constraint guidance achieves the lowest Intersection Volume, Intersection Frequency and FID than other baselines, which shows that our method achieves the lowest level of penetration while ensuring the highest generation quality. In addition, other baseline methods using our physical constraint guidance significantly reduce the degree of penetration, which demonstrates the effectiveness and universality of our guidance method. Although our guidance method effectively suppresses penetration, it also leads to a slight increase in FID, as FID only measures the similarity between generated results and the ground truth distribution and the dataset itself exists a certain degree of penetration.
4.4 Ablation Study
Network Prediction. As depicted in Sec. 3.1, a straightforward and effective strategy is to estimate clean body poses directly through a neural network, i.e., $\mathbf{x}_{1}$-prediction. We compared it with $\mathbf{v}$-prediction and the results are listed on the Prediction setting in Tab. 5 and  Tab. 6. Obviously, $\mathbf{x}_{1}$-prediction has demonstrated superior performance across both online and offline settings. The reason we analyze it is that the geometric losses to regularize the generative network during the training phase directly acts on the predicted clean body poses, while $\mathbf{v}$-prediction requires using the predicted vector field to estimate the clean poses through Equation 13, so the models trained by $\mathbf{x}_{1}$-prediction will achieve stronger overall performance in the task of human action-reaction synthesis.
Physical constraint guidance. As we discussed earlier, the results in Tab. 4 indicate that our physical constraint guidance can effectively suppress penetration occurrences, and it may also lead to a decrease in some other metrics shown in Tab. 5 and Tab. 6. We also provide a qualitative comparison of the effects before and after using our physical constraint guidance in Fig. 10.1. The qualitative and quantitative results demonstrated that our method achieves the lowest penetration level while maintaining the best quality of generated reactions.
Number of Euler sampling timesteps. We present comprehensive evaluation results in both online and offline scenarios, with varying Euler sampling intervals (2, 5, 10 and 100 timesteps), including the latency of reaction generation per frame on online settings and overall latency on offline settings. The experimental results, as detailed in Tab. 5 and Tab. 6, suggest that the 5-timestep Euler sampling consistently achieves optimal performance, demonstrating superior FID scores while maintaining low latency across both evaluation settings. Thus, we adopt the 5-timestep inference as the standard configuration like  [69] for all the experimental results reported in this study.
5 Conclusion
In this work, we have presented Action-Reaction Flow Matching (ARFlow), a novel framework for human action-reaction synthesis that addresses the limitations of existing diffusion-based approaches. By establishing direct action-to-reaction mappings through flow matching, ARFlow eliminates the need for complex conditional mechanisms and computationally expensive noise-to-data transformations. ARFlow involves a $\mathbf{x}_{1}$-prediction sampling algorithm with physical constraint guidance to enable efficient, physically plausible motion generation while preventing body penetration artifacts. Extensive evaluations on the NTU120 and Chi3D datasets demonstrate that ARFlow excels over existing methods, showing superior performance in terms of Fréchet Inception Distance and motion diversity. Additionally, it significantly reduces body collisions, as evidenced by our new Intersection Volume and Intersection Frequency metrics.
However, while ARFlow’s physical constraint guidance proves effective in reducing body penetrations, it may lead to manifold distortions—deviations from the natural motion distribution established by flow matching. Addressing this challenge opens a promising avenue for future research, focusing on developing guidance mechanisms that ensure physical plausibility without compromising motion authenticity.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No.62406195, No.62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University) and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
Supplementary Material
6 Algorithm derivation
We denote the deterministic functions: $\hat{\mathbf{x}}_{1}=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]$ as the $\mathbf{x}_{1}$-prediction, $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=u_{t}(\mathbf{x}_{t})$ as the $\mathbf{v}$-prediction. By defining the conditional probability path as a linear interpolation between $p_{0}$ and $p_{1}$, the intermediate process becomes:
$\mathbf{x}_{t}=t\mathbf{x}_{1}+[1-(1-\sigma_{\text{min}})t]\mathbf{x}_{0},$ (18)
where $\sigma_{\text{min}}>0$ is a small amount of noise. Take the derivative of $t$ on both sides:
$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0},$ (19)
In the marginal velocity formula, we obtain:
$u_{t}(\mathbf{x}_{t})$ $=\mathbb{E}[\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0}|\mathbf{x}_{t},c]$
$=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},c].$ (20)
Substitute $\mathbf{x}_{0}=\frac{\mathbf{x}_{t}-t\mathbf{x}_{1}}{1-(1-\sigma_{\text{min}})t}$ from Eq. 18 into the above equation:
$~{}~{}~{}~{}~{}~{}~{}u_{t}(\mathbf{x}_{t})$ $=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\frac{\mathbf{x}_{t}-t~{}\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]}{1-(1-\sigma_{\text{min}})t}$
$=\frac{\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t}$
where we have used the fact that $E[\mathbf{x}_{t}|\mathbf{x}_{t}]=\mathbf{x}_{t}$. According to $\hat{\mathbf{x}}_{1}=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]$ , $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=u_{t}(\mathbf{x}_{t})$, we get the equivalent form of parameterization:
$\mathbf{v}_{\theta}(\mathbf{x}_{t}$ $,t,c)=\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t},$ (21)
To simplify notations, here we omit the small amount of noise $\sigma_{\text{min}}$.
$\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-\mathbf{x}_{t}}{1-t}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (22)
Substitute Eq. 22 into the following equation:
$\mathbf{x}_{t^{\prime}}$ $=\mathbf{x}_{t}-(t-t^{\prime})\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)$
$=\mathbf{x}_{t}-(t-t^{\prime})\frac{\hat{\mathbf{x}}_{1}-\mathbf{x}_{t}}{1-t}$
$=\frac{(1-t)\mathbf{x}_{t}+(t-t^{\prime})\mathbf{x}_{t}-(t-t^{\prime})\hat{\mathbf{x}}_{1}}{1-t}$
$=\frac{1-t^{\prime}}{1-t}\mathbf{x}_{t}+\frac{t^{\prime}-t}{1-t}\hat{\mathbf{x}}_{1}.$ (23)
Finally, let $t=t_{n}$ and $t^{\prime}=t_{n+1}$, we can obtain the estimation of $\hat{\mathbf{x}}_{1}$ from Eq. 22:
$\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$
and our sampling formulation based on $\mathbf{x}_{1}$-prediction from Eq. 23:
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1}.$
Algorithm 2 Euler Sampling algorithm with physical constraint guidance. ($\mathbf{v}$-prediction)
1:  Input: $\mathcal{L}_{\text{pene}}$ the loss function ; $\mathbf{v}$ and $\theta$ the vector field predictor with pretrained parameters
2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the scale factor to control the strength of guidance
3:  Sample $\mathbf{x}_{0}$ from the action distribution
4:  for $n=1,2,...,N-1$ do
5:     # Estimate $\hat{\mathbf{x}}_{1}$ (Eq. 13)
6:     $\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
7:     # Flow maching $\mathbf{v}$-prediction sampling (Eq. 11)
8:     $\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\mathbf{x}_{t_{n}}+(t_{n+1}-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
9:     # Physical constraint guidance (Eq. 12)
10:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
11:  end for
12:  Return: The reaction motion after guidance $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
7 Parameter analysis of guidance strength
We conduct a parameter analysis of guidance strength in Tab. 7.1. The experiment results show that as the guiding strength increases, the degree of penetration between actors and reactors decreases significantly, while FID increases slightly. This is because the ground truth itself has a certain degree of penetration. Thus, this task requires our new metrics and FID to collaborate in evaluating the quality of the generated results. When the guidance strength increases to a certain extent, the decrease in penetration degree is no longer significant. Therefore, we ultimately choose $\lambda_{\text{pene}}=10$. Our method achieves the lowest penetration level while maintaining the best generation quality.
Table 7.1: Parameter analysis of guidance strength on the online, unconstrained setting on NTU120-AS. Bold indicates the best result.
Parameter settings FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
Real 0.09 38.38% 3.33
     $\lambda_{\text{pene}}$ 0 7.89 17.44% 1.55
1 8.29 16.01% 1.43
2 8.51 13.30% 1.15
5 8.65 9.60% 0.81
10 8.75 8.56% 0.76
15 9.04 8.50% 0.75
8 Details of our framework
We present our Human Action-Reaction Flow Matching (ARFlow) framework, illustrated in Fig. 2, which comprises a flow module and a Transformer decoder $G$. Given a paired action-reaction sequence and an optional signal $c$ (e.g. , an action label, dotted lines in Fig. 2), $<$$x_{0}^{1:H}$, $x_{1}^{1:H}$, $c$$>$, $x_{1}^{1:H}$ represents the reaction to generate. For a sampled timestep $t$, we linearly interpolate $x_{1}^{1:H}$ and $x_{0}^{1:H}$ as Eq. 4 to produce the $x^{1:H}_{t}$. Then the $x^{1:H}_{t}$ turns into the latent features through an FC layer to dimension d. The timestep $t$ and the optional condition $c$ are separately projected to dimension $d$ using feed-forward networks and combined to form the token $z$. The Transformer decoder $G$, implemented with stacked 8 layers, prevents future information leakage through masked multi-head attention, enabling online generation as in [69]. Decoder $G$ takes $z$ as input tokens and $x^{1:H}_{t}$ combined with a standard positional embedding as output tokens, along with a directional attention mask to ensure the model cannot access future actions at the current timestep. The decoder’s output is projected back to produce the predicted clean body poses $\hat{x}_{1}^{1:H}$. Online reaction generation is achieved in an auto-regressive manner, following the approach of [69]. The intention branch can be activated when the actor’s intention is accessible to the reactor, or deactivated otherwise. The directional attention mask can be turned off for offline settings.
At the inference stage, we employ our physical constraint guidance method. After training latent linear layers and Transformer decoder $G$, our ARFlow uses them for $x_{1}$-prediction based sampling. The sampling process is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
Figure 8.1: Human perceptual study results on NTU120-AS.
Figure 8.2: Human perceptual study results on Chi3D-AS.
9 Implementation details
Our ARFlow model is trained with $T=1,000$ timesteps using a classifier-free approach [23]. The number of decoder layers is 8 and the latent dimension of the Transformer tokens is 512. The batch size is configured as 32 for NTU120-AS and 16 for Chi3D-AS. The interaction loss weight is set to $\lambda_{\text{inter}}=1$. Each model is trained for 500K steps on single NVIDIA 4090 GPU within 48 hours. During inference, unless otherwise stated, we employ 5-timestep sampling for all the diffusion-based and our models in our experiments as [69] for a fair comparison. For the physical constraint guidance, we set the safe distance $\zeta=1$ and $\lambda_{\text{pene}}=10$
10 Details of the metric calculations
We follow the prior works in human action-reaction synthesis, ReGenNet [69] and MDM [59] to calculate the Frechet Inception Distance(FID) [22], action recognition accuracy, diversity and multi-modality. For a fair comparison, we use the pre-trained action recognition model in [69], which is a slightly modified version of ST-GCN [70]. The model takes the 6D rotation representation of the SMPL-X parameters as input and outputs classification results of action-reaction pairs. We generate 1,000 reaction samples by sampling actor motions from test sets and evaluate each method 20 times using different random seeds to calculate the average with the 95% confidence interval.
1) Frechet Inception Distance (FID) [22] measures the similarity in feature space between predicted and ground-truth motion; 2) Action Recognition Accuracy (Acc.) assesses how likely a generated motion can be successfully recognized. We adopt the pre-trained ST-GCN model to classify the generated results; 3) Diversity (Div.) evaluates feature diversity within generated motions. Given the motion feature vectors of generated motions and real motions as {$v_{1},\cdots,v_{S_{d}}$} and {$v_{1}^{\prime},\cdots,v_{S_{d}}^{\prime}$}, the diversity is defined as $Diversity=\frac{1}{S_{d}}\sum_{i=1}^{S_{d}}||v_{i}-v_{i}^{\prime}||_{2}$. $S_{d}=200$ in our experiments. 4) Multi-modality (Multimod.) quantifies the ability to generate multiple different motions for the same action type. Given a collection of motions containing $C$ action types, for $c$-th action, we randomly sample two subsets of size $S_{l}$, and then extract the corresponding feature vectors as {$v_{c,1},\cdots,v_{c,S_{l}}$} and {$v_{c,1}^{\prime},\cdots,v_{c,S_{l}}^{\prime}$}, the multimodality is defined as $Multimod.=\frac{1}{C\times S_{l}}\sum_{c=1}^{C}\sum_{i=1}^{S_{l}}||v_{c,i}-v_{c,i}^{\prime}||_{2}$. $S_{l}=20$ in our experiments. 5) Intersection Volume (IV) measures the volume of human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both. 6) Intersection Frequency (IF) measures the frequency of inter-penetration. We generate 260 samples for evaluation.
Figure 10.1: Visualization comparison of the effects before and after using physical constraint guidance. Blue for actors and Green for reactors.
11 Extra qualitative results
We show the generated motions of our method against others in Fig. 11.2. We highlight the implausible motions in rectangle marks, it is clear that our method learns the correct reactions and avoids human-human inter-penetrations as much as possible.
Figure 11.1: Failure case of our method.
Figure 11.2: The extra qualitative experiment. Blue for actors and Green for reactors.
Failure case. We also show the failure cases of our motion generation pipeline in Fig. 11.1. Our model cannot guarantee absolute physical authenticity, for example, ensuring that the hand contacts but does not penetrate during handshaking. Incorporating more sophisticated physical constraints may solve the failure cases and be considered in future.
12 User Study
We conducted a human perceptual study to investigate the quality of the motions generated by our model. We invite 20 users to provide four comparisons. For each comparison, we ask the users “Which of the two motions is more realistic?”, and each user is provided 10 sequences to evaluate.
The results are shown in Fig. 8.1 and Fig. 8.2. Our results were preferred over the other state-of-the-art and are even competitive with ground truth motions.

### v3

@@
 To address these issues, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that establishes direct action-to-reaction mappings, eliminating the need for complex conditional mechanisms.
-Our approach introduces two key innovations: an $\mathbf{x}_{1}$-prediction method that directly outputs human motions instead of velocity fields, enabling explicit constraint enforcement; and a training-free, gradient-based physical guidance mechanism that effectively prevents body penetration artifacts during sampling.
-Extensive experiments on NTU120 and Chi3D datasets demonstrate that ARFlow not only outperforms existing methods in terms of Fréchet Inception Distance and motion diversity but also significantly reduces body collisions, as measured by our new Intersection Volume and Intersection Frequency metrics.
-Figure 1: Our proposed Human Action-Reaction Flow (ARFlow).
-We directly establish a mapping between the action and reaction distribution and our sampling process is further guided by our physical constraint guidance.
-The change of colors represents the variation of the $h$-frame reaction with sampling timestep $t_{n}$.
-$\dagger$\$$\dagger$\$footnotetext: Corresponding authors.
+Our approach introduces a physical guidance mechanism specifically designed for Flow Matching (FM) that effectively prevents body penetration artifacts during sampling.
+Moreover, we discover the bias of traditional flow matching sampling algorithm and employ a reprojection method to revise the sampling direction of FM.
+To further enhance the reaction diversity, we incorporate randomness into the sampling process.
+Extensive experiments on NTU120, Chi3D and InterHuman datasets demonstrate that ARFlow not only outperforms existing methods in terms of Fréchet Inception Distance and motion diversity but also significantly reduces body collisions, as measured by our new Intersection Volume and Intersection Frequency metrics.
+Our project is available at https://arflow2025.github.io/.
 1 Introduction
@@
 ARFlow establishes direct pathways between action and reaction distributions, enabling simpler training and faster inference compared to diffusion-based methods.
-Instead of traditional vector field prediction ($\mathbf{v}$-prediction) in standard flow matching [38, 25], ARFlow directly outputs clean body poses ($\mathbf{x}_{1}$-prediction).
-This innovation allows direct application of geometric constraints during training and better alignment with motion generation objectives.
-We further develop a specialized sampling algorithm optimized for $\mathbf{x}_{1}$-prediction in human motion generation.
-To eliminate unrealistic body collisions between characters, we introduce a penetration-aware sampling process that maintains physical plausibility through gradient guidance without compromising motion quality.
+Instead of traditional vector field prediction ($\mathbf{v}$-prediction) in standard flow matching [40, 27], ARFlow directly outputs clean body poses ($\mathbf{x}_{1}$-prediction) in human motion generation.
+To eliminate unrealistic body collisions between characters, we further develop a specialized sampling algorithm that performs physical constraint guidance at $\mathbf{x}_{1}$ to avoid inaccurate gradients, and projects it back onto a revised Flow Matching path through interpolation.
+This innovation maintains physical plausibility through gradient guidance without compromising motion quality.
 Additionally, we introduce two new physics-aware evaluation metrics: Intersection Volume (IV) and Intersection Frequency (IF).
@@
   • We propose ARFlow, the first Flow Matching architecture that creates direct pathways between human action and reaction distributions, eliminating the need for complex conditional mechanisms prevalent in diffusion approaches.
-  • We propose a new $\mathbf{x}_{1}$-prediction method under flow matching specifically designed to integrate physical constraints during both training and sampling.
-This combined approach successfully prevents body collisions between characters while preserving natural movement.
+  • We integrate physical constraints and employ a reprojection method to revise the sampling direction of FM, surpassing vanilla guidance methods.
+This successfully prevents body collisions between characters while preserving natural movement.
+To further enhance the reaction diversity, we introduce randomness into the sampling process.
   • We introduce two new evaluation metrics (IV, IF) for interaction quality assessment.
 Experiments demonstrate state-of-the-art performance in Fréchet Inception Distance (FID), MultiModality (Multimod.), IV and IF on NTU120 and Chi3D datasets.
-Figure 2: Pipeline of ARFlow.
-(a) For a sampled timestep $t$, we linearly interpolate a coupled action-reaction pair as Eq.
-4 to produce the intermediate state $x^{1:H}_{t}$, which is then turns into a d-dimensional latent feature through a linear layer.
-We use Transformer Decoder Units to directly predict clean reaction motions.
-(b) After training the networks in (a), our ARFlow uses them for $\mathbf{x}_{1}$-$prediction$ based sampling.
-The sampling process is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
+Figure 1: Our proposed Human Action-Reaction Flow (ARFlow).
+We directly establish a mapping between the action and reaction distribution and our sampling process is further guided by our physical constraint guidance.
+The change of colors represents the variation of the $h$-frame reaction with sampling timestep $t_{n}$.
 2 Related Work
@@
 However, it often produce physically-implausible inter-penetrations between the actor and reactor since they disregard physical constraints in the generative process.
-Our method addresses this problem by guiding the sampling process with physical constraint guidance.
+Our method addresses this problem by ARFlow sampling with physical constraint guidance.
 2.3 Flow Matching
-Flow Matching [38, 42, 1, 45] has emerged as an efficient alternative to diffusion models, offering linear generation trajectories through ODE solvers.
+Flow Matching [40, 44, 1, 48, 47, 17] has emerged as an efficient alternative to diffusion models, offering linear generation trajectories through ODE solvers.
 This paradigm enables simplified training and accelerated inference [16, 41], with successful applications spanning images [40, 16], audio [36], video [5], and point clouds [68].
@@
 3.2.
+Figure 2: Pipeline of ARFlow.
+(a) For a sampled timestep $t$, we linearly interpolate a coupled action-reaction pair as Eq.
+4 to produce the intermediate state $x^{1:H}_{t}$, which is then turns into a d-dimensional latent feature through a linear layer.
+We use Transformer Decoder Units to directly predict clean reaction motions.
+(b) After training the networks in (a), our ARFlow uses them for sampling, which is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
 3.1 Human Action-Reaction Flow Matching
@@
 The training objective of our flow model is as follows:
-${\mathcal{L}_{\text{fm}}=\mathbb{E}_{\mathbf{x}_{1}\sim q(\mathbf{x}_{1}),\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),t\sim[0,1]}[\|\mathbf{x}_{1}-G_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}].}$ (5)
+${~{}~{}~{}~{}\mathcal{L}_{\text{fm}}=\mathbb{E}_{\mathbf{x}_{1}\sim q(\mathbf{x}_{1}),\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),t\sim[0,1]}[\|\mathbf{x}_{1}-G_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}].}$ (5)
 Following [72], we employ explicit interaction losses to evaluate the relative distances of body pose $\bm{\theta}(\mathbf{x}_{1},\mathbf{x}_{0})$, orientation $\bm{q}(\mathbf{x}_{1},\mathbf{x}_{0})$ and translation $\bm{\gamma}(\mathbf{x}_{1},\mathbf{x}_{0})$ between the actor and reactor.
@@
 Since our neural network outputs $\hat{\mathbf{x}}_{1}$, we require to construct an equivalent relationship between the neural network’s predictions of $\mathbf{v}$ and $\mathbf{x}_{1}$.
-The equivalent form of parameterization Equation 21 derived from our appendix is as follows:
-${\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t},}$ (7)
-Then, our sampling based on $\mathbf{x}_{1}$-prediction can be achieved by first sampling $\mathbf{x}_{0}$ and then solving Equation 1 employing an ODE solver [52, 33, 2] through our trained neural network $\hat{\mathbf{x}}_{1}=G_{\theta}(\mathbf{x}_{t},t,c)$.
+The equivalent form of parameterization Eq.
+23 derived from our appendix is as follows:
+${\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ (7)
+Then, our sampling based on $\mathbf{x}_{1}$-prediction can be achieved by first sampling $\mathbf{x}_{0}$ and then solving Eq.
+1 employing an ODE solver [55, 35, 2] through our trained neural network $\hat{\mathbf{x}}_{1}=G_{\theta}(\mathbf{x}_{t},t,c)$.
 We use the Euler ODE solver and discretization process involves dividing the procedure into $N$ steps, leading to the following formulation:
@@
 where the integer time step $t_{1}=0<t_{2}<\cdots<t_{N}=1$.
-By using equivalent form of parameterization Equation 7, we finally obtain our flow maching sampling formulation based on $\mathbf{x}_{1}$-prediction:
-$\mathbf{x}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1},$ (9)
-which is more suitable for human motion generation. Detailed derivation is provided in the supplementary.
-3.2 Physical Constraint Guidance
-There exist many physically implausible inter-penetrations between the actor and reactor in the generated results of current diffusion-based methods [69, 59, 15], owing to the stochastic sampling process that disregards physical constraints.
-To address the issue, we propose a training-free physical constraint guidance method that calculates the penetration gradients $\nabla\mathcal{L}_{\text{pene}}$ that pulls the reactor in the opposite direction to the actor.
-Then these gradients are employed to guide the sampling process at each timestep to ensure the suppression of penetration.
-Specifically, We use $\psi^{h}_{i}(\mathbf{x})$ to represent the position of the $i$-th joint in the $h$-th frame of the generated reaction motion $\mathbf{x}$.
-Inspired by [30, 36], we calculate the signed distance function (SDF) between the actor and the reactor.
-Thus, our penetration loss functon is defined as:
-$\mathcal{L}_{\text{pene}}(\mathbf{x})\vcentcolon=\sum_{i,h}-\mathrm{min}\Bigl{(}\mathrm{SDF}(\psi^{h}_{i}(\mathbf{x})),\zeta\Bigr{)}$ (10)
-where $\zeta$ defines the safe distance between the actor and the reactor, beyond which the gradient becomes zero, and $\mathrm{SDF}$ is the signed distance function for an actor in the $h$-th frame, which dynamically changes across frames.
-Flow maching sampling with physical constraint guidance.
-At each sampling step, the gradient of $\nabla\mathcal{L}_{\text{pene}}$ is utilized to guide the generative process.
-In contrast to the standard process for motion synthesis in Equation 8, the sampling process with physical constraint guidance is formalized as:
-$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\mathbf{x}_{t_{n}}+(t_{n+1}-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (11)
-$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (12)
-where $\mathbf{x}^{\prime}_{t_{n+1}}$ denoted the intermediate results of standard sampling in Equation 8.
-However, standard sampling based on $\mathbf{v}$-prediction often requires to estimate $\hat{\mathbf{x}}_{1}$:
-$\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (13)
-from current state $\mathbf{x}_{t_{n}}$ using an equivalent form of parameterization in Equation 7 to input into the signed distance function $\mathcal{L}_{\text{pene}}(\mathbf{x})$.
-In comparison, the sampling process based on $\mathbf{x}_{1}$-prediction with physical constraint guidance is as follows:
-$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1},$ (14)
-$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (15)
-where $\hat{x}_{1}$ is the clean body poses predicted by a neural network $G_{\theta}$, which is more compatible with our physical constraint guidance method.
-The complete Euler sampling with physical constraint guidance for $\mathbf{x}_{1}$-prediction or $\mathbf{v}$-prediction is shown in Algorithm 1 and Algorithm 2 respectively.
+By using equivalent form of parameterization Eq.
+7, we finally obtain our flow maching sampling formulation based on $\mathbf{x}_{1}$-prediction:
+$\mathbf{x}_{t_{n+1}}\leftarrow\frac{1-(1-\sigma_{\text{min}})t_{n+1}}{1-(1-\sigma_{\text{min}})t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-(1-\sigma_{\text{min}})t_{n}}~{}\hat{\mathbf{x}}_{1},$ (9)
+which is more suitable for human motion generation. Detailed derivation is provided in Appendix A.
+3.2 ARFlow Sampling with Physical Constraint Guidance
+There exist many physically implausible inter-penetrations between the actor and reactor in the generated results of current diffusion-based methods [72, 62, 15], owing to the sampling process that disregards physical constraints.
+To address the issue, we employ a penetration gradients $\nabla\mathcal{L}_{\text{pene}}$ to guide the sampling process.
+We follow [32, 38] to calculate the signed distance function (SDF) between the actor and the reactor as our penetration loss functon:
+$\mathcal{L}_{\text{pene}}(\mathbf{x})\vcentcolon=\sum_{i,h}-\mathrm{min}\Bigl{(}\mathrm{SDF}(\psi^{h}_{i}(\mathbf{x})),\zeta\Bigr{)},$ (10)
+where $\psi^{h}_{i}(\mathbf{x})$ represents the position of the $i$-th joint in the $h$-th frame of the generated reaction motion $\mathbf{x}$, the $\zeta$ defines the safe distance between the actor and the reactor, beyond which the gradient becomes zero, and $\mathrm{SDF}$ is the signed distance function for an actor in the $h$-th frame, which dynamically changes across frames.
+Traditional diffusion guidance methods [12, 31, 63] first estimate $\hat{\mathbf{x}}_{1}$ from current state $\mathbf{x}_{t_{n}}$ with a denoiser network $\epsilon_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$, and then calculate gradients of the loss function with respect to current state $\mathbf{x}_{t_{n}}$, so it inevitably requires differentiation of the neural network, resulting in inaccurate gradients $\nabla\mathcal{L}_{\text{pene}}$.
+To avoid this issue, we directly updates the gradient at $\hat{\mathbf{x}}_{1}$:
+$\hat{\mathbf{x}}^{\prime}_{1}\leftarrow\hat{\mathbf{x}}_{1}-\lambda_{\text{pene}}\nabla_{\hat{\mathbf{x}}_{1}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (11)
+where $\hat{\mathbf{x}}_{1}$ is the clean body poses predicted by our neural network $G_{\theta}$ and $\lambda_{\text{pene}}$ is the guidance strength.
+Then, we use the linear interpolation of Flow Matching to project back to the intermediate state of learned FM path.
+Interpolation between action and reaction.
+Since we build flow matching between the action and reaction distribution, unlike the process of adding noise back in diffusion, our interpolation process depends on the initial point $\mathbf{x}_{0}$, so we need to project towards the direction of the initial point.
+The Flow Matching sampling algorithm Eq.
+9 is equivalent to the following formulation:
+$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{\mathbf{x}}_{0}\leftarrow~{}$ $\hat{\mathbf{x}}_{1}+\frac{\mathbf{x}_{t_{n}}-(1+\sigma_{\text{min}}t)\hat{\mathbf{x}}_{1}}{1-(1-\sigma_{\text{min}})t_{n}},$ (12)
+$\mathbf{x}_{t_{n+1}}\leftarrow~{}$ $t_{n+1}\hat{\mathbf{x}}_{1}+[1-(1-\sigma_{\text{min}})t_{n+1}]~{}\hat{\mathbf{x}}_{0}.$ (13)
+This sampling process essentially finds a $\hat{\mathbf{x}}_{0}$ along the opposite direction of the current velocity field for linear interpolation as Eq.
+12 (black dotted lines in Fig.
+3).
+Obviously, this predicted $\hat{\mathbf{x}}_{0}$ deviates from the initial point $\mathbf{x}_{0}$ (purple dotted lines in Fig.
+3).
+In order to correct the projection direction of traditional Flow Matching, we use a weight factor $w$ to weight $\hat{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}$:
+$\hat{\mathbf{x}}^{*}_{0}\leftarrow w\hat{\mathbf{x}}_{0}+(1-w)\mathbf{x}_{0}.$ (14)
+and use $\hat{\mathbf{x}}^{*}_{0}$ as our final endpoint for interpolation.
+The ARFlow sampling and traditional sampling algorithm for $\mathbf{x}_{1}$-prediction with physical constraint guidance are shown in Algorithm 2 and 1 respectively.
 In practice, we use $\mathbf{x}_{1}$-prediction for its better performance.
 Under iterative sampling and physical constraint guidance, our method can generate more realistic and physically-plausible reaction motions.
-Our guidance method is actually a refined fine-tuning, which may be not suitable for training phases.
-In addition, the loss during the training phase mainly measures the difference between generated results and ground truth, while our guidance during the inference phase can provide more flexible guidance based on the quality of the generated results.
-Metrics.
-To qualitatively measure the degree of penetration, we introduced two metrics: 1) Intersection Volume (IV).
+Stochastic sampling to enhance diversity of reactions.
+To further enhance the diversity of reactions generated by ARFlow, we incorporate randomness into the sampling process.
+The interpolation Eq.
+13 can be written in the following equivalent form:
+$\mathbf{x}_{t_{n+1}}\leftarrow\hat{\mathbf{x}}_{1}+(1-t_{n+1})(\hat{\mathbf{x}}_{0}-\hat{\mathbf{x}}_{1})+\sigma_{\text{min}}t_{n+1}\hat{\mathbf{x}}_{0}.$ (15)
+The interpolation process can be understood as a projection in the opposite direction of the current learned velocity field $\hat{\mathbf{x}}_{1}-{\hat{\mathbf{x}}}_{0}$.
+Thus, we can weight the projection direction $\hat{\mathbf{x}}^{*}_{0}-\hat{\mathbf{x}}^{\prime}_{1}$ and stochastic direction $d_{\text{random}}$ to incorporate randomness:
+$d_{\text{mix}}\leftarrow\hat{\mathbf{x}}^{*}_{0}-\hat{\mathbf{x}}^{\prime}_{1}+\beta[d_{\text{random}}-(\hat{\mathbf{x}}^{*}_{0}-\hat{\mathbf{x}}^{\prime}_{1})],$ (16)
+$\mathbf{x}_{t_{n+1}}\leftarrow~{}\hat{\mathbf{x}}^{\prime}_{1}+(1-t_{n+1})d_{\text{mix}}+\sigma_{\text{min}}t_{n+1}\hat{\mathbf{x}}^{*}_{0},$ (17)
+where $\beta$ is the factor to control the strength of randomness.
+Our guidance method is actually a refined fine-tuning, which may be not suitable for training.
+In addition, the loss during the training mainly measures the difference between generated results and ground truth, while our guidance during the inference phase can provide more flexible guidance based on the quality of the generated results.
+Figure 3: Illustration of our ARFlow sampling with physical constraint guidance.
+FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
+Real 0.09 21.96% 5.35
+       Offline
+MDM 22.04 15.45% 3.36
+ReGenNet 6.19 19.14% 3.07
+ARFlow 5.00 9.67% 1.94
+MDM ${}^{\text{vanilla}}$ 20.42 13.29% 2.26
+ReGenNet ${}^{\text{vanilla}}$ 7.69 17.85% 2.16
+[…]
+ARFlow ${}^{\text{improved}}$ 5.05 3.73% 0.23
+       Online
+MDM 51.11 32.63% 17.97
+ReGenNet 11.00 13.84% 3.50
+ARFlow 7.89 8.39% 3.26
+MDM ${}^{\text{vanilla}}$ 40.03 22.63% 10.35
+ReGenNet ${}^{\text{vanilla}}$ 11.63 13.05% 2.71
+ARFlow ${}^{\text{vanilla}}$ 8.61 3.29% 1.44
+ARFlow ${}^{\text{improved}}$ 8.07 3.23% 0.53
+Table 1: Human action-reaction synthesis with physical constraint guidance on NTU120-AS. Bold indicates the best result.
+Metrics. To qualitatively measure the degree of penetration, we introduced two metrics:
+1) Intersection Volume (IV).
 Penetrate in [76, 23] just measures ground penetration which is not suitable for measuring the degree of penetration between humans.
 Interpenetration in [46] can only be computed as rigid bodies in the physics simulation.
-Inspired by Solid Intersection Volume (IV) [78, 41] in hand-object interaction, we measure human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both.
+Inspired by Solid Intersection Volume (IV) [81, 43], we measure human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both.
 Intersection Volume (IV) is defined as
 $IV=\frac{1}{H\cdot N_{\text{total}}}\sum_{i=1}^{N_{\text{total}}}\sum_{h=1}^{H}V^{h}_{\text{pene}},$ (18)
-where $V^{h}_{\text{pene}}$ represents the intersection volume of frame $h$ and $N_{\text{total}}$ denotes the total number of samples.
-2) Intersection Frequency (IF).
-However, solely considering this metric can be misleading since it does not account for the case where there is no contact between the reactor and the reactor at all.
-Inspired by Contact Frequency in [36, 54], we introduce IF to measure the frequency of inter-penetration, i.e.
-$IF=\frac{f_{\text{pene}}}{F_{\text{total}}}\cdot 100\%,$ (17)
-where $f_{\text{pene}}$ represents the number of frames with inter-penetrations and $F_{\text{total}}$ denotes the total number of frames.
-Algorithm 1 Euler Sampling algorithm with physical constraint guidance. ($\mathbf{x}_{1}$-prediction)
+where $V^{h}_{\text{pene}}$ represents intersection volume of frame $h$ and $N_{\text{total}}$ denotes the total number of samples.
+2) Intersection Frequency (IF). Inspired by Contact Frequency in [38, 57], we introduce IF to measure the frequency of inter-penetration, i.e.
+$IF=f_{\text{pene}}/F_{\text{total}},$ (19)
+where $f_{\text{pene}}$ represents the number of inter-penetration frames and $F_{\text{total}}$ is the total number of frames.
+Algorithm 1 Sampling algorithm with vanilla guidance of physical constraints.
 1:  Input: $\mathcal{L}_{\text{pene}}$ the loss function ; $G$ and $\theta$ the clean body poses predictor with pretrained parameters
-2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the scale factor to control the strength of guidance
+2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the guidance strength
 3:  Sample $\mathbf{x}_{0}$ from the action distribution
@@
 5:     $\hat{\mathbf{x}}_{1}\leftarrow G_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
-6:     # Flow maching $\mathbf{x}_{1}$-prediction sampling (Eq. 14)
+6:     # Flow Matching $\mathbf{x}_{1}$-prediction sampling (Eq. 9)
 7:     $\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1}$
-8:     # Physical constraint guidance (Eq. 15)
-9:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
+8:     # Physical constraint guidance
+9:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}-\lambda_{\text{pene}}\nabla_{\hat{\mathbf{x}}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
 10:  end for
-11:  Return: The reaction motion after guidance $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
+11:  Return: The reaction motion $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
+Algorithm 2 Sampling algorithm with improved guidance of physical constraints.
+1:  Input: $\mathcal{L}_{\text{pene}}$ the loss function ; $G$ and $\theta$ the clean body poses predictor with pretrained parameters
+2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the guidance strength; $w$ weight factor
+3:  Sample $\mathbf{x}_{0}$ from the action distribution
+4:  for $n=1,2,...,N-1$ do
+5:     $\hat{\mathbf{x}}_{1}\leftarrow G_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
+6:     $\hat{\mathbf{x}}_{0}\leftarrow\hat{\mathbf{x}}_{1}+\frac{\mathbf{x}_{t_{n}}-(1+\sigma_{\text{min}}t)\hat{\mathbf{x}}_{1}}{1-(1-\sigma_{\text{min}})t_{n}}$       #  (Eq. 12)
+7:     # Physical constraint guidance at $\hat{\mathbf{x}}_{1}$(Eq. 11)
+8:     $\hat{\mathbf{x}}^{\prime}_{1}\leftarrow\hat{\mathbf{x}}_{1}-\lambda_{\text{pene}}\nabla_{\hat{\mathbf{x}}_{1}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
+9:     # Direction correction
+10:     $\hat{\mathbf{x}}^{*}_{0}\leftarrow w\hat{\mathbf{x}}_{0}+(1-w)\mathbf{x}_{0}$
+11:     # Interpolation (Eq. 13)
+12:     $\mathbf{x}_{t_{n+1}}\leftarrow t_{n+1}\hat{\mathbf{x}}^{\prime}_{1}+[1-(1-\sigma_{\text{min}})t_{n+1}]~{}\hat{\mathbf{x}}^{*}_{0}$
+13:  end for
+14:  Return: The reaction motion $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
 4 Experiments
 Figure 4: Qualitative comparisons of human action-reaction synthesis results. Blue for actors and Green for reactors.
-Our experiment setting of human action-reaction synthesis is online, unconstrained as in  [69] for its significant potential for practical applications.
+Our experiment setting of human action-reaction synthesis is online and unconstrained as in  [72] for its significant potential for practical applications.
 Online represents real-time reaction generation where future motions of the actor are not visible to the reactor, and the opposite is offline to relax the synchronicity.
@@
 MDM-GRU [62] $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
-ReGenNet $\underline{11.00^{\pm 0.74}}$ $\mathbf{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
+ReGenNet [72] $\underline{11.00^{\pm 0.74}}$ $\mathbf{0.749^{\pm 0.0002}}$ $13.80^{\pm 0.16}$ $\underline{22.90^{\pm 0.14}}$
 ARFlow $\mathbf{7.89^{\pm 0.18}}$ $\underline{0.743^{\pm 0.0002}}$ $\underline{13.60^{\pm 0.10}}$ $\mathbf{24.11^{\pm 0.13}}$
@@
 MDM-GRU [62] $18.63^{\pm 25.87}$ $0.574^{\pm 0.0046}$ $6.20^{\pm 0.24}$ $10.49^{\pm 0.32}$
-ReGenNet $\underline{13.76^{\pm 4.78}}$ $0.601^{\pm 0.0040}$ $6.35^{\pm 0.24}$ $\underline{12.02^{\pm 0.33}}$
+ReGenNet [72] $\underline{13.76^{\pm 4.78}}$ $0.601^{\pm 0.0040}$ $6.35^{\pm 0.24}$ $\underline{12.02^{\pm 0.33}}$
 ARFlow $\mathbf{10.92^{\pm 3.70}}$ $0.600^{\pm 0.0040}$ $\underline{6.68^{\pm 0.25}}$ $\mathbf{12.74^{\pm 0.17}}$
@@
 […]
-Real $0.09^{\pm 0.00}$ $0.867^{\pm 0.0002}$ $13.06^{\pm 0.09}$ $25.03^{\pm 0.23}$
-cVAE [32] $74.73^{\pm 4.86}$ $0.760^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.40^{\pm 0.26}$
-AGRoL [15] $16.55^{\pm 1.41}$ $0.716^{\pm 0.0002}$ $13.84^{\pm 0.10}$ $21.73^{\pm 0.20}$
-MDM [59] $7.49^{\pm 0.62}$ $\mathbf{0.775^{\pm 0.0003}}$ $\underline{13.67^{\pm 0.18}}$ $24.14^{\pm 0.29}$
-MDM-GRU [59] $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
-ReGenNet $\underline{6.19^{\pm 0.33}}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\underline{25.21^{\pm 0.34}}$
-ARFlow $\mathbf{5.00^{\pm 0.17}}$ $\underline{0.772^{\pm 0.0002}}$ $13.84^{\pm 0.09}$ $\mathbf{25.10^{\pm 0.17}}$
-Table 3: Results on the offline, unconstrained setting on NTU120-AS. We highlight the best result in Bold and the second best in underline.
-Table 4: Human action-reaction synthesis with physical constraint guidance on NTU120-AS. Bold indicates the best result.
-FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
-Real 0.09 38.38% 3.33
-       Offline
-MDM 22.04 29.77% 2.50
-ReGenNet 6.19 35.77% 2.85
-ARFlow 5.00 23.91% 2.01
-MDM ${}^{\text{guidance}}$ 20.42 25.58% 2.10
-ReGenNet ${}^{\text{guidance}}$ 7.69 25.10% 1.92
-ARFlow ${}^{\text{guidance}}$ 5.19 15.82% 1.31
-       Online
-MDM 51.11 51.87% 4.22
-ReGenNet 11.00 31.92% 2.63
-ARFlow 7.89 17.44% 1.55
-MDM ${}^{\text{guidance}}$ 40.03 44.84% 3.67
-ReGenNet ${}^{\text{guidance}}$ 11.63 17.11% 1.37
-ARFlow ${}^{\text{guidance}}$ 8.75 8.56% 0.76
-[…]
 Real $0.085^{\pm 0.0003}$ $0.867^{\pm 0.0002}$ $13.063^{\pm 0.0908}$ $25.032^{\pm 0.2332}$
@@
 2) $v$ $14.726^{\pm 0.2143}$ $0.743^{\pm 0.0002}$ $14.154^{\pm 0.0923}$ $23.329^{\pm 0.1125}$
-Guidance w. $\mathcal{L}_{\text{pene}}$ $8.751^{\pm 0.2047}$ $0.740^{\pm 0.0002}$ $13.613^{\pm 0.1004}$ $24.096^{\pm 0.1433}$
+Guidance w. $\mathcal{L}_{\text{pene}}$ $8.073^{\pm 0.1981}$ $0.741^{\pm 0.0002}$ $13.713^{\pm 0.1079}$ $24.077^{\pm 0.1370}$
 Timesteps 2 $15.965^{\pm 0.2728}$ $0.733^{\pm 0.0002}$ $13.740^{\pm 0.0896}$ $26.767^{\pm 0.1440}$ $0.055$
@@
 Table 4: Ablation studies on the online, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
-[…]
-Real $0.085^{\pm 0.0003}$ $0.867^{\pm 0.0002}$ $13.063^{\pm 0.0908}$ $25.032^{\pm 0.2332}$
-Prediction 1) $x_{1}$ $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$
-2) $v$ $7.585^{\pm 0.1562}$ $0.757^{\pm 0.0002}$ $13.775^{\pm 0.0982}$ $24.200^{\pm 0.1355}$
-Guidance w. $\mathcal{L}_{\text{pene}}$ $5.196^{\pm 0.1144}$ $0.753^{\pm 0.0002}$ $13.824^{\pm 0.0839}$ $25.164^{\pm 0.1630}$
-Timesteps 2 $7.936^{\pm 0.1581}$ $0.759^{\pm 0.0002}$ $14.538^{\pm 0.1016}$ $25.904^{\pm 0.1754}$ $0.023$
-5 $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$ $0.053$
-10 $5.506^{\pm 0.1657}$ $0.744^{\pm 0.0002}$ $13.870^{\pm 0.0942}$ $24.732^{\pm 0.1533}$ $0.110$
-100 $5.836^{\pm 0.3763}$ $0.748^{\pm 0.0002}$ $13.635^{\pm 0.0948}$ $24.058^{\pm 0.1371}$ $1.132$
-Best ARFlow $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$ $0.053$
-Table 6: Ablation studies on the offline, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
 4.1 Experiment setup
@@
 Datasets.
-We evaluate our model on NTU120-AS and Chi3D-AS datasets with SMPL-X [46] body models and actor-reactor annotations as in [47].They contains 8118 and 373 human interaction sequences, respectively.
-“AS”[69] represents that they are an extended version of the original dataset [17, 40, 61], which adds annotations to distinguish actor-reactor order of each interaction sequence and SMPL-X body models for more detailed representations.
+We evaluate our model on NTU120-AS, Chi3D-AS and InterHuman-AS datasets with SMPL-X [49] body models and actor-reactor annotations as in [50].They contains 8118, 373 and 6022 human interaction sequences, respectively.
+“AS”[72] represents that they are an extended version of the original dataset [18, 42, 64, 39], which adds annotations to distinguish actor-reactor order of each interaction sequence and SMPL-X body models for more detailed representations.
 We adopt the 6D rotation representation [82] in all our experiments.
@@
 2 and Chi3D-AS dataset in Tab.
-2, our proposed ARFlow notably outperforms baselines in terms of the FID metric, demonstrating that our method better models the mapping between the action and the reaction distribution.
+3, our proposed ARFlow notably outperforms baselines in terms of the FID metric, demonstrating that our method better models the mapping between the action and the reaction distribution.
 Our method achieves the best FID and multi-modality, second best for the action recognition accuracy and diversity on the NTU120-AS dataset and the best FID and multi-modality, second best for the diversity on the Chi3D-AS dataset.
@@
 Given the restricted size of the Chi3D-AS test set, some fluctuations in the experimental results are to be expected.
+The results of InterHuman-AS dataset in Appendix B.1 show our method also yields the best results compared to baselines.
 As for the offline setting, we replace the Transformer decoder units equipped with attention masks with an 8-layer Transformer encoder architecture just like ReGenNet [72].
 As demonstrated in Tab.
-3, our model achieves superior performance on most of the metrics, highlighting the effectiveness and versatility of our method across various scenarios.
+B.2, our model also achieves superior performance on most of the metrics.
 4.3 Physical Constraint Guidance
@@
 In Fig.
-3, we provide a qualitative comparison of reaction sequences generated by MDM, ReGenNet and our ARFlow with physical constraint guidance, for the same action sequences sampled from the test sets of NTU120-AS datasets.
-The visualization results demonstrate that both MDM and ReGenNet produce varying degrees of penetration between the actor and the reactor.
+4, for the same action sequences sampled from the test sets of NTU120-AS datasets, visualization results demonstrate that both MDM and ReGenNet produce varying degrees of penetration between the actor and the reactor.
 In contrast, our method not only produces more physically plausible reactions, i.e., mitigate penetration problems, but also more responsive reactions, i.e., move backward when kicked, due to stronger modeling ability for causal relationship between actions and reactions.
@@
 In Tab.
-4, our ARFlow with physical constraint guidance achieves the lowest Intersection Volume, Intersection Frequency and FID than other baselines, which shows that our method achieves the lowest level of penetration while ensuring the highest generation quality.
-In addition, other baseline methods using our physical constraint guidance significantly reduce the degree of penetration, which demonstrates the effectiveness and universality of our guidance method.
+1, our ARFlow with physical constraint guidance achieves the lowest Intersection Volume, Intersection Frequency and FID than other baselines, which shows that our method achieves the lowest level of penetration while ensuring the highest generation quality.
+In addition, the performance is further improved through our ARFlow sampling with improved physical constraint guidance, which demonstrates the effectiveness of our guidance method.
 Although our guidance method effectively suppresses penetration, it also leads to a slight increase in FID, as FID only measures the similarity between generated results and the ground truth distribution and the dataset itself exists a certain degree of penetration.
@@
 Network Prediction.
-As depicted in Sec.
+As depicted in Tab.
 3.1, a straightforward and effective strategy is to estimate clean body poses directly through a neural network, i.e., $\mathbf{x}_{1}$-prediction.
@@
 4 and  Tab.
-6.
-Obviously, $\mathbf{x}_{1}$-prediction has demonstrated superior performance across both online and offline settings.
-The reason we analyze it is that the geometric losses to regularize the generative network during the training phase directly acts on the predicted clean body poses, while $\mathbf{v}$-prediction requires using the predicted vector field to estimate the clean poses through Equation 13, so the models trained by $\mathbf{x}_{1}$-prediction will achieve stronger overall performance in the task of human action-reaction synthesis.
+B.3.
+Obviously, $\mathbf{x}_{1}$-prediction has demonstrated superior performance across both settings.
+The reason we analyze it is that the geometric losses to regularize the generative network during the training phase directly acts on the predicted clean body poses, while $\mathbf{v}$-prediction requires using the predicted vector field to estimate the clean poses, so the models trained by $\mathbf{x}_{1}$-prediction will achieve stronger overall performance.
 Physical constraint guidance.
@@
 4 and Tab.
-6.
-We also provide a qualitative comparison of the effects before and after using our physical constraint guidance in Fig.
-10.1.
+B.3.
+We also provide a qualitative comparison of the effects before and after using our physical constraint guidance in Tab.
+H.1.
 The qualitative and quantitative results demonstrated that our method achieves the lowest penetration level while maintaining the best quality of generated reactions.
@@
 4 and Tab.
-6, suggest that the 5-timestep Euler sampling consistently achieves optimal performance, demonstrating superior FID scores while maintaining low latency across both evaluation settings.
+B.3, suggest that the 5-timestep Euler sampling consistently achieves optimal performance, demonstrating superior FID scores while maintaining low latency across both evaluation settings.
 Thus, we adopt the 5-timestep inference as the standard configuration like  [72] for all the experimental results reported in this study.
@@
 By establishing direct action-to-reaction mappings through flow matching, ARFlow eliminates the need for complex conditional mechanisms and computationally expensive noise-to-data transformations.
-ARFlow involves a $\mathbf{x}_{1}$-prediction sampling algorithm with physical constraint guidance to enable efficient, physically plausible motion generation while preventing body penetration artifacts.
-Extensive evaluations on the NTU120 and Chi3D datasets demonstrate that ARFlow excels over existing methods, showing superior performance in terms of Fréchet Inception Distance and motion diversity.
+ARFlow involves a reprojection sampling algorithm with physical constraint guidance to enable efficient, physically plausible motion generation while preventing body penetration artifacts.
+Extensive evaluations on the NTU120, Chi3D and InterHuman datasets demonstrate that ARFlow excels over existing methods, showing superior performance in terms of Fréchet Inception Distance and motion diversity.
 Additionally, it significantly reduces body collisions, as evidenced by our new Intersection Volume and Intersection Frequency metrics.
-However, while ARFlow’s physical constraint guidance proves effective in reducing body penetrations, it may lead to manifold distortions—deviations from the natural motion distribution established by flow matching.
-Addressing this challenge opens a promising avenue for future research, focusing on developing guidance mechanisms that ensure physical plausibility without compromising motion authenticity.
+Limitations.
+Although we attempt to use a reprojection method to address the issue of manifold distortions—deviations from the natural motion distribution established by flow matching, this problem still exists and our penetration loss function is just a simple design.
+Addressing these challenge opens a promising avenue for future theoretical research, focusing on developing guidance mechanisms that ensure physical plausibility without compromising motion authenticity.
 Acknowledgement
-This work was supported by National Natural Science Foundation of China (No.62406195, No.62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University) and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
-Supplementary Material
-6 Algorithm derivation
+This work was supported by National Natural Science Foundation of China (No.62406195, No.62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University) and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
+Supplementary Materials
+[…]
+[…]
+Appendix
+Appendix A Algorithm derivation
 We denote the deterministic functions: $\hat{\mathbf{x}}_{1}=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]$ as the $\mathbf{x}_{1}$-prediction, $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=u_{t}(\mathbf{x}_{t})$ as the $\mathbf{v}$-prediction.
 By defining the conditional probability path as a linear interpolation between $p_{0}$ and $p_{1}$, the intermediate process becomes:
-$\mathbf{x}_{t}=t\mathbf{x}_{1}+[1-(1-\sigma_{\text{min}})t]\mathbf{x}_{0},$ (18)
+$\mathbf{x}_{t}=t\mathbf{x}_{1}+[1-(1-\sigma_{\text{min}})t]\mathbf{x}_{0},~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (20)
 where $\sigma_{\text{min}}>0$ is a small amount of noise. Take the derivative of $t$ on both sides:
-$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0},$ (19)
+$\frac{d\mathbf{x}_{t}}{dt}=\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0},~{}~{}~{}~{}~{}~{}~{}$ (21)
 In the marginal velocity formula, we obtain:
 $u_{t}(\mathbf{x}_{t})$ $=\mathbb{E}[\mathbf{x}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{0}|\mathbf{x}_{t},c]$
-$=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},c].$ (20)
+$=\mathbb{E}[\mathbf{x}_{1}|\mathbf{x}_{t},c]-(1-\sigma_{\text{min}})\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},c].~{}~{}~{}~{}$ (22)
 Substitute $\mathbf{x}_{0}=\frac{\mathbf{x}_{t}-t\mathbf{x}_{1}}{1-(1-\sigma_{\text{min}})t}$ from Eq. 20 into the above equation:
@@
 $\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)$ $=\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t},$ (23)
-To simplify notations, here we omit the small amount of noise $\sigma_{\text{min}}$.
-$\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-\mathbf{x}_{t}}{1-t}.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (22)
 Substitute Eq. 23 into the following equation:
-$\mathbf{x}_{t^{\prime}}$ $=\mathbf{x}_{t}-(t-t^{\prime})\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)$
-$=\mathbf{x}_{t}-(t-t^{\prime})\frac{\hat{\mathbf{x}}_{1}-\mathbf{x}_{t}}{1-t}$
-$=\frac{(1-t)\mathbf{x}_{t}+(t-t^{\prime})\mathbf{x}_{t}-(t-t^{\prime})\hat{\mathbf{x}}_{1}}{1-t}$
-$=\frac{1-t^{\prime}}{1-t}\mathbf{x}_{t}+\frac{t^{\prime}-t}{1-t}\hat{\mathbf{x}}_{1}.$ (23)
+$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathbf{x}_{t^{\prime}}$ $=\mathbf{x}_{t}-(t-t^{\prime})\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)$
+$=\mathbf{x}_{t}-(t-t^{\prime})\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t}$
+$=\frac{1-(1-\sigma_{\text{min}})t^{\prime}}{1-(1-\sigma_{\text{min}})t}\mathbf{x}_{t}+\frac{t^{\prime}-t}{1-(1-\sigma_{\text{min}})t}\hat{\mathbf{x}}_{1}.$ (24)
 Finally, let $t=t_{n}$ and $t^{\prime}=t_{n+1}$, we can obtain the estimation of $\hat{\mathbf{x}}_{1}$ from Eq. 23:
-$\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$
+$\hat{\mathbf{x}}_{1}\leftarrow(1-\sigma_{\text{min}})\mathbf{x}_{t_{n}}+(1-(1-\sigma_{\text{min}})t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (25)
 and our sampling formulation based on $\mathbf{x}_{1}$-prediction from Eq. 24:
-$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1}.$
-Algorithm 2 Euler Sampling algorithm with physical constraint guidance. ($\mathbf{v}$-prediction)
+$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-(1-\sigma_{\text{min}})t_{n+1}}{1-(1-\sigma_{\text{min}})t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-(1-\sigma_{\text{min}})t_{n}}~{}\hat{\mathbf{x}}_{1}.$ (26)
+Algorithm 3 Sampling algorithm with vanilla guidance of physical constraints. ($\mathbf{v}$-prediction)
 1:  Input: $\mathcal{L}_{\text{pene}}$ the loss function ; $\mathbf{v}$ and $\theta$ the vector field predictor with pretrained parameters
@@
 5:     # Estimate $\hat{\mathbf{x}}_{1}$ (Eq. 25)
-6:     $\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
+6:     $\hat{\mathbf{x}}_{1}\leftarrow(1-\sigma_{\text{min}})\mathbf{x}_{t_{n}}+(1-(1-\sigma_{\text{min}})t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
 7:     # Flow maching $\mathbf{v}$-prediction sampling (Eq. 8)
 8:     $\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\mathbf{x}_{t_{n}}+(t_{n+1}-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$
-9:     # Physical constraint guidance (Eq. 12)
+9:     # Physical constraint guidance
 10:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
@@
 12:  Return: The reaction motion after guidance $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
-7 Parameter analysis of guidance strength
+Appendix B Experimental results on the InterHuman-AS dataset and offline settings
+For the text-conditioned setting, we adopt T2M [20], MDM [62], MDM-GRU [62], RAIG [60] and InterGen [39] as baselines.
+Tab.
+B.1 shows our method also yields the best results compared to baselines.
+To demonstrate the universality of our ARFlow and physical guidance methods, we also conducted offline setting experiments in Tab.B.2 and Tab.B.3.
+[…]
+Real $0.722^{\pm 0.004}$ $0.002^{\pm 0.0002}$ $3.503^{\pm 0.011}$ $5.390^{\pm 0.058}$
+T2M [20] $0.224^{\pm 0.003}$ $32.482^{\pm 0.0975}$ $7.299^{\pm 0.016}$ $4.350^{\pm 0.073}$ $0.719^{\pm 0.041}$
+MDM [62] $0.370^{\pm 0.006}$ $3.397^{\pm 0.0352}$ $8.640^{\pm 0.065}$ $4.780^{\pm 0.117}$ $2.288^{\pm 0.039}$
+MDM-GRU [62] $0.328^{\pm 0.012}$ $6.397^{\pm 0.2140}$ $8.884^{\pm 0.040}$ $4.851^{\pm 0.081}$ $2.076^{\pm 0.040}$
+RAIG [60] $0.363^{\pm 0.008}$ $2.915^{\pm 0.0292}$ $7.294^{\pm 0.027}$ $4.736^{\pm 0.099}$ $2.203^{\pm 0.049}$
+InterGen [39] $0.374^{\pm 0.005}$ $13.237^{\pm 0.0352}$ $10.929^{\pm 0.026}$ $4.376^{\pm 0.042}$ $\mathbf{2.793^{\pm 0.014}}$
+ReGenNet [72] $0.407^{\pm 0.003}$ $2.265^{\pm 0.0969}$ $6.860^{\pm 0.0040}$ $5.214^{\pm 0.139}$ $2.391^{\pm 0.023}$
+ARFlow $\mathbf{0.434^{\pm 0.003}}$ $\mathbf{1.637^{\pm 0.0413}}$ $\mathbf{3.949^{\pm 0.0042}}$ $\mathbf{5.259^{\pm 0.117}}$ $2.502^{\pm 0.021}$
+Table B.1: Comparison to state-of-the-arts on the online, unconstrained setting for human action-reaction synthesis on the InterHuman-AS dataset.
+$\rightarrow$ denotes that the result closer to Real is better, and $\pm$ represents 95% confidence interval.
+We highlight the best result in Bold.
+[…]
+Real $0.09^{\pm 0.00}$ $0.867^{\pm 0.0002}$ $13.06^{\pm 0.09}$ $25.03^{\pm 0.23}$
+cVAE [34] $74.73^{\pm 4.86}$ $0.760^{\pm 0.0002}$ $11.14^{\pm 0.04}$ $18.40^{\pm 0.26}$
+AGRoL [15] $16.55^{\pm 1.41}$ $0.716^{\pm 0.0002}$ $13.84^{\pm 0.10}$ $21.73^{\pm 0.20}$
+MDM [62] $7.49^{\pm 0.62}$ $\mathbf{0.775^{\pm 0.0003}}$ $\underline{13.67^{\pm 0.18}}$ $24.14^{\pm 0.29}$
+MDM-GRU [62] $24.25^{\pm 1.39}$ $0.720^{\pm 0.0002}$ $\mathbf{13.43^{\pm 0.09}}$ $22.24^{\pm 0.29}$
+ReGenNet [72] $\underline{6.19^{\pm 0.33}}$ $0.772^{\pm 0.0003}$ $14.03^{\pm 0.09}$ $\underline{25.21^{\pm 0.34}}$
+ARFlow $\mathbf{5.00^{\pm 0.17}}$ $\underline{0.772^{\pm 0.0002}}$ $13.84^{\pm 0.09}$ $\mathbf{25.10^{\pm 0.17}}$
+Table B.2: Results on the offline, unconstrained setting on NTU120-AS. We highlight the best result in Bold and the second best in underline.
+[…]
+Real $0.085^{\pm 0.0003}$ $0.867^{\pm 0.0002}$ $13.063^{\pm 0.0908}$ $25.032^{\pm 0.2332}$
+Prediction 1) $x_{1}$ $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$
+2) $v$ $7.585^{\pm 0.1562}$ $0.757^{\pm 0.0002}$ $13.775^{\pm 0.0982}$ $24.200^{\pm 0.1355}$
+Guidance w. $\mathcal{L}_{\text{pene}}$ $5.048^{\pm 0.1167}$ $0.750^{\pm 0.0002}$ $13.838^{\pm 0.0893}$ $25.048^{\pm 0.1595}$
+Timesteps 2 $7.936^{\pm 0.1581}$ $0.759^{\pm 0.0002}$ $14.538^{\pm 0.1016}$ $25.904^{\pm 0.1754}$ $0.023$
+5 $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$ $0.053$
+10 $5.506^{\pm 0.1657}$ $0.744^{\pm 0.0002}$ $13.870^{\pm 0.0942}$ $24.732^{\pm 0.1533}$ $0.110$
+100 $5.836^{\pm 0.3763}$ $0.748^{\pm 0.0002}$ $13.635^{\pm 0.0948}$ $24.058^{\pm 0.1371}$ $1.132$
+Best ARFlow $\mathbf{5.003^{\pm 0.1654}}$ $\mathbf{0.762^{\pm 0.0002}}$ $13.844^{\pm 0.0905}$ $\mathbf{25.104^{\pm 0.1704}}$ $0.053$
+Table B.3: Ablation studies on the offline, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
+Appendix C Influence of sampling randomness
+As depicted in Tab. C.1, although stochastic sampling increases the diversity of generated reaction motions, it sometimes has some impact on the quality of the sample due to its stochastic nature.
+[…]
+Real $0.085^{\pm 0.0003}$ $0.867^{\pm 0.0002}$ $13.063^{\pm 0.0908}$ $25.032^{\pm 0.2332}$
+Randomness $\beta$ 0.05 $13.821^{\pm 0.2895}$ $0.709^{\pm 0.0003}$ $14.002^{\pm 0.1055}$ $24.269^{\pm 0.1363}$
+0.02 $8.060^{\pm 0.1517}$ $0.728^{\pm 0.0002}$ $13.928^{\pm 0.1076}$ $24.161^{\pm 0.1512}$
+0.01 $7.671^{\pm 0.1357}$ $0.728^{\pm 0.0002}$ $13.895^{\pm 0.1080}$ $24.114^{\pm 0.1486}$
+ARFlow 0 $7.894^{\pm 0.1814}$ $\mathbf{0.743^{\pm 0.0002}}$ $\mathbf{13.599^{\pm 0.1005}}$ $24.105^{\pm 0.1310}$
+Table C.1: Randomness Influence studies on the online, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
+Appendix D Parameter analysis of guidance strength and weight factor
 We conduct a parameter analysis of guidance strength in Tab.
-7.1.
-The experiment results show that as the guiding strength increases, the degree of penetration between actors and reactors decreases significantly, while FID increases slightly.
+D.1.
+The result of the experiments show that as the guiding strength increases, the degree of penetration between actors and reactors decreases significantly, while FID increases slightly.
 This is because the ground truth itself has a certain degree of penetration.
@@
 When the guidance strength increases to a certain extent, the decrease in penetration degree is no longer significant.
-Therefore, we ultimately choose $\lambda_{\text{pene}}=10$.
+Therefore, we ultimately choose $\lambda_{\text{pene}}=2$.
 Our method achieves the lowest penetration level while maintaining the best generation quality.
-Table 7.1: Parameter analysis of guidance strength on the online, unconstrained setting on NTU120-AS. Bold indicates the best result.
+As for the weight factor, the results show that the minimum value of FID does not occur at the endpoints, thus demonstrating the effectiveness of our weighting method.
+Table D.1: Parameter analysis of guidance strength and weight factor on the online, unconstrained setting on NTU120-AS. Bold indicates the best result.
 Parameter settings FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
-Real 0.09 38.38% 3.33
-     $\lambda_{\text{pene}}$ 0 7.89 17.44% 1.55
-1 8.29 16.01% 1.43
-2 8.51 13.30% 1.15
-5 8.65 9.60% 0.81
-10 8.75 8.56% 0.76
-15 9.04 8.50% 0.75
-8 Details of our framework
+Real 0.09 21.96% 5.35
+     $\lambda_{\text{pene}}$ 0 7.89 8.39% 3.26
+1 7.98 5.80% 1.15
+2 8.20 3.54% 0.68
+5 8.49 1.22% 0.13
+10 9.41 0.78% 0.21
+     $w$ 0 8.37 2.71% 0.35
+0.1 8.30 2.78% 0.36
+0.3 8.19 2.96% 0.37
+0.5 8.11 3.12% 0.41
+0.7 8.07 3.23% 0.53
+0.9 8.08 3.32% 0.64
+1 8.20 3.54% 0.68
+Appendix E Details of our framework
 We present our Human Action-Reaction Flow Matching (ARFlow) framework, illustrated in Fig.
@@
 The sampling process is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
-Figure 8.1: Human perceptual study results on NTU120-AS.
-Figure 8.2: Human perceptual study results on Chi3D-AS.
-9 Implementation details
+Appendix F Implementation details
 Our ARFlow model is trained with $T=1,000$ timesteps using a classifier-free approach [25].
 The number of decoder layers is 8 and the latent dimension of the Transformer tokens is 512.
-The batch size is configured as 32 for NTU120-AS and 16 for Chi3D-AS.
+The batch size is configured as 32 for NTU120-AS, InterHuman-AS and 16 for Chi3D-AS.
 The interaction loss weight is set to $\lambda_{\text{inter}}=1$.
@@
 During inference, unless otherwise stated, we employ 5-timestep sampling for all the diffusion-based and our models in our experiments as [72] for a fair comparison.
-For the physical constraint guidance, we set the safe distance $\zeta=1$ and $\lambda_{\text{pene}}=10$
-10 Details of the metric calculations
+For the physical constraint guidance, we set the safe distance $\zeta=0.5$, $\lambda_{\text{pene}}=2$ and $w=0.7$
+Figure F.1: Human perceptual study results on NTU120-AS.
+Figure F.2: Human perceptual study results on Chi3D-AS.
+Appendix G Details of the metric calculations
 We follow the prior works in human action-reaction synthesis, ReGenNet [72] and MDM [62] to calculate the Frechet Inception Distance(FID) [24], action recognition accuracy, diversity and multi-modality.
@@
 We generate 260 samples for evaluation.
-Figure 10.1: Visualization comparison of the effects before and after using physical constraint guidance. Blue for actors and Green for reactors.
-11 Extra qualitative results
-We show the generated motions of our method against others in Fig.
-11.2.
-We highlight the implausible motions in rectangle marks, it is clear that our method learns the correct reactions and avoids human-human inter-penetrations as much as possible.
-Figure 11.1: Failure case of our method.
-Figure 11.2: The extra qualitative experiment. Blue for actors and Green for reactors.
-Failure case.
-We also show the failure cases of our motion generation pipeline in Fig.
-11.1.
-Our model cannot guarantee absolute physical authenticity, for example, ensuring that the hand contacts but does not penetrate during handshaking.
-Incorporating more sophisticated physical constraints may solve the failure cases and be considered in future.
-12 User Study
+Appendix H User Study
 We conducted a human perceptual study to investigate the quality of the motions generated by our model.
@@
 For each comparison, we ask the users “Which of the two motions is more realistic?”, and each user is provided 10 sequences to evaluate.
-The results are shown in Fig. 8.1 and Fig. 8.2. Our results were preferred over the other state-of-the-art and are even competitive with ground truth motions.
+The results are shown in Fig. F.1 and Fig. F.2. Our results were preferred over the other state-of-the-art and are even competitive with ground truth motions.
+Figure H.1: Visualization comparison of the effects before and after using physical constraint guidance. Blue for actors and Green for reactors.
+Appendix I Extra qualitative results
+We show the generated motions of our method against others in Fig.
+I.1.
+We highlight the implausible motions in rectangle marks, it is clear that our method learns the correct reactions and avoids human-human inter-penetrations as much as possible.
+Figure I.1: The extra qualitative experiment. Blue for actors and Green for reactors.
+Figure I.2: Failure case of our method.
+Failure case.
+We also show the failure cases of our motion generation pipeline in Fig.
+I.2.
+Our model cannot guarantee absolute physical authenticity, for example, ensuring that the hand contacts but does not penetrate during handshaking.
+Incorporating more sophisticated physical constraints may solve the failure cases and be considered in future.
+Appendix J Broader Impacts
+Our model demonstrates significant potential for AR/VR and gaming applications by enabling the generation of plausible human reactions.
+Beyond virtual environments, the proposed approach provides an innovative technical pathway for real-world human-robot interaction, where motion patterns can be transferred to robotic systems through motion remapping technology.
+Although this advancement may inspire future research, we acknowledge potential misuse risks similar to other generative models, warranting ethical considerations as the technology develops.